OpenStudy (anonymous):

Can you show me how to calculate: (16y^4)^3/4

OpenStudy (anonymous):

start with \[16=2^4\] so you have \[(2^4y^4)^{\frac{3}{4}}\]

OpenStudy (turingtest):

I assume you mean (16y^4)^(3/4) remember (ax^n)^m=a^m(x^nm) so in this case (16y^4)^(3/4)=16^(3/4)(y^3)=2^3(y^3)=8y^3

OpenStudy (anonymous):

then multiply the exponents, which really means cancel the 4s and you get \[2^3y^3=8y^3\]

OpenStudy (anonymous):